OpenStudy (anonymous):

4 A man weighs 750 N on the surface of the earth. What would be his weight when standing on the moon? The masses of the earth and the moon are respectively, 5.98×1024 kg and 7.36×1022 kg. Their radii are respectively 6.37×103 km and 1.74×103 km

Parth (parthkohli):

Weight on any planet is the same as the gravitational force exerted by it. Gravitational force is given by\[G\dfrac{M_{\rm planet}m_{\rm you}}{R^2}\]

Parth (parthkohli):

So, it's given that your weight is 750 N on Earth. That means that the Earth exerts a gravitational force of 750 N. Let's first find your mass.\[\rm 750 N = \dfrac{G \times M_{\large earth}\times m_{you}}{(R_{\large earth})^2} \]

Parth (parthkohli):

Once you find your mass, plug it into the expression for the gravitational force exerted by the moon on you. That'd be your weight on the moon.